If x-1/3y=3, then the value of 3x-y-6=?
Answer by Maricel Balos(8)

Sol'n:
x-1/3y=3 then 3x-y-6=?
x-1=3(3y)
x-1=9y
x=9y+1
3x-y-6=3(9y+1)
3x-y-6=27y+3
3x=27y+y+3+6
3x=28y+9
x=28y/3+9/3
28y/3+9/3=9y+1
28y+9=3(9y+1)
28y+9=27y+3
28y-27y=3-9
y=-6
x=9y+1
x=9(-6)+1
x=-54+1
x=-53
therefore the sol'n is x=-53 and y=-6 the sol'n set is (-53,-6)
the graph is a straight line that lies on the 4th quadrant